OpenStudy (anonymous):

1:what is the simplified form of x^4-81/x+3? 2:what is the simplified form of (x^2yz)^2(xy^2z^2)/(xyz)^2

OpenStudy (zzr0ck3r):

\[x^4-81=(x^2)^2-(3^2)^2=(x^2-3^2)(x^2+9)=(x-3)(x+3)(x^2+9)\]

OpenStudy (zzr0ck3r):

does that help?